If you take a 2.0 m long simple pendulum to Pluto, how would the period of motion compare to when on Earth?
Answer #1

L = 2m

Since gravity on Pluto is gp = g/12

Now T = 2.pai √L/g

=) T = 2×3.143×√(2/9.8) = 2.84s on earth

But On Pluto Tp = 2×3.143×√(12×2/9.8) = 9.837s

That means T(Pluto) = √12 × T(earth) = (3.46) time that on earth .

Thank you. Please give a thumb up.